OpenStudy (anonymous):

You weigh a sample of potassium hydrogen phthalate (KHP) to prepare a primary standard for a titration. You later discover that the KHP was contaminated with sugar. To determine the amount of KHP in the mixture, you take 6.336 g of the mixture and make a 100.0 mL solution. You then titrate 10.00 mL of this solution with a 0.1787 M sodium hydroxide solution. You find that 13.75 mL of the NaOH solution is needed to reach the endpoint. What is the percent by mass of KHP in the mixture?

OpenStudy (anonymous):

we have Veq = 13.75 ml of (0.1787 M) we do : C1V1 = C2Veq ===> C1 = (C2Veq)/V1 AN : C1 = (0.1787 * 13.75)/ 10 = 0.2457 M so in 1000ml we have 0.2457 mol of KPH but we worked on 100ml so n = 0.02457mol === > m(KPH)= n*M(KPH) = 0.02457 * M finally 6.336 g ----> 100% and m(KPH) -----> X % ................done ........... sorry about my english writing .......

OpenStudy (anonymous):

@ljones69